OpenStudy (anonymous):

Solve the equation: 9x − 2 = 4(x + 3)

OpenStudy (anonymous):

9x-2=4x+12 =>9x-4x=12+2 =>5x=14 =>x=14/5

OpenStudy (anonymous):

thankyou *****